OpenStudy (anonymous):

(Medal for best answer and explanation) Solve 7x = 2x – 35

OpenStudy (anonymous):

Isolate the variable by subtracting from each side: 5x=-35 Divide by 5 to get x=5

OpenStudy (anonymous):

Thanks!

OpenStudy (merchandize):

7x-2x=-35 5x=-35 x=-5 like terms should be in either sides so that you can subtract and divide easily.

OpenStudy (anonymous):

ok, So how exactly would I solve this next one. 1/4(20-4a)= 6- a

OpenStudy (anonymous):

Are they all real numbers?

OpenStudy (merchandize):

first of all remove the fraction then solve and find the value of a

OpenStudy (anonymous):

(20-4a)=6- a So then we subtract 4 from the 20 which leaves us with 16a right? But then what happens to the other side?

OpenStudy (merchandize):

first of all to remove the fraction we have to think for such digit by which we can remove the left hand side denominator part i.e,remove 1/4

OpenStudy (merchandize):

1/4(20-4a)=6-a multiply both side by 4. 4*1/4(20-4a)=4(6-a) 20-4a=24-4a 20-24=-4a+4a 20-24=0 so it's is a false statement.

OpenStudy (merchandize):

have you got that??

OpenStudy (anonymous):

Ohhh.. okay. I get it now, c: Thank you so much

OpenStudy (anonymous):

x = -7

OpenStudy (merchandize):

oh that's a blunder....it's really x=-7

OpenStudy (merchandize):

thank you shay.

OpenStudy (anonymous):

@EnvyLeague answer is -7 don't forget : )

OpenStudy (usman1995):

yes absolutely -35/5=-7 ,, answer is -7

OpenStudy (mathstudent55):

try this: \(7x = 2x – 35\) \(5x = -35\) \(x = -7\)

OpenStudy (mathstudent55):

Distribute the 1/4 on the left side since both 20 and 4 are divisible by 4. 1/4(20-4a)= 6- a 5 - a = 6 - a Add a to both sides: 5 = 6 Answer: no solution. The equation is false.